# Investigating charmed hybrid baryons via QCD sum rules

### v1

@@ ACKNOWLEDGMENTS
Figure 4: Variations of $M_{+}$ as functions of the threshold value $\omega_{c}$ and the Borel mass $T$. In the left panel the long-dashed, solid, and short-dashed curves correspond to fixed Borel masses $T=0.37~{}\mathrm{GeV}$, $0.38~{}\mathrm{GeV}$, and $0.40~{}\mathrm{GeV}$, respectively. In the right panel the long-dashed, solid, and short-dashed curves correspond to fixed threshold values $\omega_{c}=2.85~{}\mathrm{GeV}$, $2.95~{}\mathrm{GeV}$, and $3.05~{}\mathrm{GeV}$, respectively. These results are obtained using the current $J_{\Sigma_{cg}^{+},\bar{3}c,1,1,1/2^{+}}$, which is associated with the doublet $[\Sigma_{cg}^{+},\bar{3}_{c},1,1,G]$.
Similarly, we apply the QCD sum rule method to study the multiplets $[\bar{3}_{f},\bar{3}_{c},0,1,G]$, $[6_{f},\bar{3}_{c},1,1,G]$, $[6_{f},\bar{3}_{c},1,2,G]$, and $[6_{f},\bar{3}_{c},1,0,G]$. The obtained results are summarized in Table. 1. Additionally, we have examined the charmed hybrid baryons belonging to the other four multiplets—$[6_{f},6_{c},0,1,G]$, $[\bar{3}_{f},6_{c},1,0,G]$, $[\bar{3}_{f},6_{c},1,1,G]$, and $[\bar{3}_{f},6_{c},1,2,G]$—as well as those with the negative parity. We find that they have significantly larger masses, primarily because the relevant two-point correlation functions are negative and thus non-physical when $\omega$ is not sufficiently large.

### v2

@@ IV Results at the ${\mathcal{O}}(1/m_{Q})$ order
Figure 5: Variations of $M_{+}$ as functions of (a) the threshold value $\omega_{c}$ and (b) the Borel mass $T$. In the left panel the long-dashed, solid, and short-dashed curves correspond to fixed Borel masses $T=0.37~\mathrm{GeV}$, $0.38~\mathrm{GeV}$, and $0.40~\mathrm{GeV}$, respectively. In the right panel the long-dashed, solid, and short-dashed curves correspond to fixed threshold values $\omega_{c}=2.85~\mathrm{GeV}$, $2.95~\mathrm{GeV}$, and $3.05~\mathrm{GeV}$, respectively. These results are obtained using the current $J_{\Sigma_{cg}^{+},\bar{3}c,1,1,1/2^{+}}$, which is associated with the doublet $\Sigma_{cg}^{+},\bar{3}_{c},1,1,G]$.
Similarly, we apply the QCD sum rule method to study the other currents belonging to the multiplets $\bar{3}_{f},\bar{3}_{c},0,1,G]$, $6_{f},\bar{3}_{c},1,1,G]$, $6_{f},\bar{3}_{c},1,2,G]$, and $6_{f},\bar{3}_{c},1,0,G]$. The obtained results are summarized in Table 1, where the current $J_{6_{f},\bar{3}_{c},1,1,1/2^{+}}$ belonging to the doublet $6_{f},\bar{3}_{c},1,1,G]$ couples to the lowest-lying charmed hybrid baryon. A point of confusion arises because this current exhibits a symmetric flavor structure for the two light quarks, which contrasts with the flavor structure of the lowest-lying charmed baryon $\Lambda_{cg}$, but is consistent with the flavor structure of $\Sigma_{cg}$. To explain the possible reasons, we notice that there are two states with the flavor structure $\bar{3}_{f}$, but five with $6_{f}$. Hence, if the mass of one of the five states with $6_{f}$ were to decrease, it would become possible for this state to become the lowest-lying one, leading to the situation where the lowest-lying charmed baryons are $\Sigma_{cg}$ rather than $\Lambda_{cg}$. Indeed, one can observe in Table 1 that only one $\Sigma_{cg}$ state is lower than the $\Lambda_{cg}(1/2^{+})$ state belonging to the $\bar{3}_{f},\bar{3}_{c},0,1,G]$ doublet, while all the others are higher than this $\Lambda_{cg}(1/2^{+})$ state.
[…]
Before concluding this section, we would like to note that in this calculation, we have focused on the ${\mathcal{O}}(1/m_{Q})$ corrections, particularly those related to the propagators in the QCD sum rules, as described above. However, the ${\mathcal{O}}(1/m_{Q})$ corrections also affect the interpolating currents, known as the flow corrections, whose effects are not considered in the present study. Although the flow corrections are of the same order, their effects are generally smaller and mainly affect the fine details of the coupling constants and Dirac structures. Therefore, they are expected to have a relatively smaller effect compared to the propagator corrections Neubert:1993mb ; Grozin:1997ih . For simplicity, we have neglected these corrections in the current analysis, but they will be systematically explored in future studies.
Table 1: Parameters of the charmed hybrid baryons belonging to the multiplets $\bar{3}_{f},\bar{3}_{c},0,1,G]$, $6_{f},\bar{3}_{c},1,0,G]$, $6_{f},\bar{3}_{c},1,1,G]$, and $6_{f},\bar{3}_{c},1,2,G]$, calculated using the QCD sum rule method within the framework of heavy quark effective theory.
(GeV) (GeV) (GeV) (GeV) (GeV) (GeV4)
$\bar{3}_{f},\bar{3}_{c},0,1,G]$ 3.20 $0.400\leq T\leq 0.410$ $2.72^{+0.17}_{-0.15}$ $\Lambda_{cg}(1/2^{+})$ $4.74^{+0.23}_{-0.20}$ $0.62^{+0.12}_{-0.10}$ $0.155^{+0.031}_{-0.024}$
$\Lambda_{cg}(3/2^{+})$ $5.36^{+0.30}_{-0.26}$ $0.045^{+0.009}_{-0.007}$
3.45 $0.409\leq T\leq 0.437$ $2.76^{+0.11}_{-0.09}$ $\Xi_{cg}(1/2^{+})$ $4.91^{+0.15}_{-0.14}$ $0.54^{+0.10}_{-0.08}$ $0.176^{+0.023}_{-0.019}$
$\Xi_{cg}(3/2^{+})$ $5.45^{+0.20}_{-0.17}$ $0.051^{+0.007}_{-0.005}$
$6_{f},\bar{3}_{c},1,0,G]$ $2.75$ $0.394\leq T\leq 0.396$ $2.25\pm 0.12$ $\Sigma_{cg}(1/2^{+})$ $4.75^{+0.21}_{-0.20}$ $0.091^{+0.015}_{-0.013}$
3.00 $0.373\leq T\leq 0.420$ $2.30^{+0.14}_{-0.07}$ $\Xi^{\prime}_{cg}(1/2^{+})$ $4.88^{+0.20}_{-0.13}$ $0.104^{+0.021}_{-0.011}$
3.25 $0.357\leq T\leq 0.442$ $2.41^{+0.15}_{-0.08}$ $\Omega_{cg}(1/2^{+})$ $5.09^{+0.21}_{-0.14}$ $0.183^{+0.041}_{-0.019}$
$6_{f},\bar{3}_{c},1,1,G]$ 2.95 $0.368\leq T\leq 0.398$ $2.46^{+0.15}_{-0.14}$ $\Sigma_{cg}(1/2^{+})$ $3.36^{+0.27}_{-0.26}$ $2.16^{+0.32}_{-0.29}$ $0.079^{+0.015}_{-0.013}$
$\Sigma_{cg}(3/2^{+})$ $5.51^{+0.29}_{-0.26}$ $0.046^{+0.009}_{-0.007}$
3.20 $0.351\leq T\leq 0.423$ $2.50\pm 0.08$ $\Xi^{\prime}_{cg}(1/2^{+})$ $3.60\pm 0.20$ $2.15^{+0.30}_{-0.27}$ $0.090\pm 0.010$
$\Xi^{\prime}_{cg}(3/2^{+})$ $5.74^{+0.21}_{-0.19}$ $0.052\pm 0.006$
3.45 $0.351\leq T\leq 0.447$ $2.62\pm 0.08$ $\Omega_{cg}(1/2^{+})$ $3.82\pm 0.21$ $2.14^{+0.30}_{-0.27}$ $0.158^{+0.018}_{-0.017}$
$\Omega_{cg}(3/2^{+})$ $5.96^{+0.22}_{-0.19}$ $0.091^{+0.010}_{-0.010}$
$6_{f},\bar{3}_{c},1,2,G]$ 3.50 $0.418\leq T\leq 0.429$ $2.98^{+0.16}_{-0.14}$ $\Sigma_{cg}(3/2^{+})$ $6.63^{+0.36}_{-0.30}$ $1.39^{+0.21}_{-0.23}$ $0.364^{+0.063}_{-0.051}$
$\Sigma_{cg}(5/2^{+})$ $5.24^{+0.25}_{-0.22}$ $0.123^{+0.021}_{-0.017}$
3.75 $0.428\leq T\leq 0.457$ $3.04^{+0.10}_{-0.09}$ $\Xi^{\prime}_{cg}(3/2^{+})$ $6.76^{+0.28}_{-0.24}$ $1.35^{+0.18}_{-0.20}$ $0.422^{+0.052}_{-0.043}$
$\Xi^{\prime}_{cg}(5/2^{+})$ $5.42^{+0.17}_{-0.16}$ $0.143^{+0.018}_{-0.015}$
4.00 $0.434\leq T\leq 0.484$ $3.16\pm 0.09$ $\Omega_{cg}(3/2^{+})$ $6.99^{+0.28}_{-0.24}$ $1.31^{+0.18}_{-0.20}$ $0.740^{+0.080}_{-0.073}$
$\Omega_{cg}(5/2^{+})$ $5.69^{+0.17}_{-0.16}$ $0.250^{+0.027}_{-0.025}$
V Discussions and Conclusion
In recent years substantial theoretical and experimental efforts have been devoted to the study of glueballs and hybrid mesons. However, their internal structures remain elusive, underscoring the necessity for continued investigations—particularly into the comparatively less-explored hybrid baryons. In this paper we employ the QCD sum rule method within the framework of heavy quark effective theory to investigate charmed hybrid baryons with various quantum numbers. We systematically construct twenty-eight interpolating currents for these baryons and utilize seven of them to perform numerical analyses. These seven currents form the four baryon multiplets $\bar{3}_{f},\bar{3}_{c},0,1,G]$, $6_{f},\bar{3}_{c},1,0,G]$, $6_{f},\bar{3}_{c},1,1,G]$, and $6_{f},\bar{3}_{c},1,2,G]$. Altogether, we compute the masses of nineteen states with quark-gluon configurations qqcgqqcg, qscgqscg, and sscgsscg ($q=u/d$). The resulting mass predictions are summarized in Table 1.
In particular, the charmed hybrid baryons belonging to the doublet $6_{f},\bar{3}c,1,1,G]$ comprise six states: $\Sigma_{cg}(1/2^{+})$, $\Sigma_{cg}(3/2^{+})$, $\Xi^{\prime}_{cg}(1/2^{+})$, $\Xi^{\prime}_{cg}(3/2^{+})$, $\Omega_{cg}(1/2^{+})$, and $\Omega_{cg}(3/2^{+})$. Among these, the three spin-$1/2$ states are predicted to be the lowest in mass:
$M_{\Sigma_{cg}(1/2^{+})}$ $=$ $3.36^{+0.27}_{-0.26}~\rm{GeV}\,$,
$M_{\Xi^{\prime}_{cg}(1/2^{+})}$ $=$ $3.59\pm 0.20~\rm{GeV}\,$, (66)
$M_{\Omega_{cg}(1/2^{+})}$ $=$ $3.82\pm 0.21~\rm{GeV}\,.$
It is important to note that all these states have positive parity. Recalling the case of hybrid mesons, the MIT bag model Barnes:1982tx ; Chanowitz:1982qj suggested that the lightest supermultiplet states correspond to those with $J^{PC}=1^{--}$, $0^{-+}$, $1^{-+}$, and $2^{-+}$, all of which have negative parity, suggesting it possible for the lightest hybrid baryons to have positive parity. However, we must acknowledge that the situation for hybrid baryons is quite complex due to the interplay between quark-gluon dynamics and the QCD vacuum structure, so the presence of gluonic excitations in hybrid baryons has the potential to lead to a different parity pattern compared to the naive constituent parton model, in which gluonic degrees of freedom are not explicitly considered. These hybrid states still involve many uncertainties, and thus, extensive further research is needed to fully understand their properties.
Charmed hybrid baryons can decay via the excitation of a color-octet $\bar{q}q$ or $\bar{s}s$ pair from the valence gluon, followed by the recombination of the remaining color-octet clusters—qqcqqc, qscqsc, or sscssc—with the excited $\bar{q}q$ or $\bar{s}s$ pair into color-singlet baryons and mesons. Representative decay channels for the charmed hybrid baryons belonging to the doublet $6_{f},\bar{3}_{c},1,1,G]$ are listed in Table 2, providing viable avenues for their experimental identification in future studies.
Table 2: Possible decay channels of the charmed hybrid baryons belonging to the doublet $6_{f},\bar{3}_{c},1,1,G]$.
States Decay patterns
$\Sigma_{cg}(1/2^{+})$ NDND, $\Sigma D_{s}$, $ND^{*}$, $\Sigma D_{s}^{*}$, $\Delta D$, $\Sigma^{*}D_{s}$
$\Omega_{cg}(1/2^{+})$ $\Xi D$, $\Xi D^{*}$, $\Xi^{*}D$, $\Omega D_{s}$, $\Xi^{*}D^{*}$
ACKNOWLEDGMENTS